12.50 A small, opaque, diffuse object at T, = 400 K is sus- pended in a large furnace whose interior walls are at T; = 2000 KPLEAS SHOW ALL WORK, Explain your steps for part b pleas!
Answer #1

Given Data: To = 4ook. TE = 2000k 0.7 E = 0.20 0.5 1 3 (Mm) a) Total Emissivity and al alsoppitivity of Susface : d, Ts = 1(4And now, A, Tf = 1 (2000) -> 2000Mm.k. 12 T = 3 (2000) Good Mmok. From Hackbody radiation functions table : for 2000 Mnak, Fl6) Reflected Radiative fiux: .: Goef P. G. = (1-2) ELTA = (1-0.601) (5.67x1681d/m24) (2000k)4. = 3.62 +105 W/m² Geref Net Rad

If you like this answer then please give me Upvote.. Thank you